OpenStudy (anonymous):

Simplify. Assume that all expressions under radicals represent positive numbers.

OpenStudy (anonymous):

\[6\sqrt{y}+9\sqrt{y}\]

OpenStudy (anonymous):

If you have 6 oranges and you add 9 oranges, how many oranges will you have?

OpenStudy (anonymous):

Let \(\sqrt{y} = oranges\)